### v4

# Inducing Causal World Models in LLMs for Zero-Shot Physical Reasoning

Abstract
Large Language Models (LLMs), despite their advanced linguistic capabilities, fundamentally lack an intuitive understanding of physical dynamics, limiting their application in real-world scenarios requiring causal reasoning. This paper introduces Causal World Model Induction (CWMI), a novel framework designed to explicitly embed a model of causal physics within an LLM. Our approach integrates a specialized Causal Physics Module (CPM) and a new training objective, the Causal Intervention Loss ($L_{causal}$), which forces the model to learn cause-and-effect relationships from multimodal data. By training the model to predict the outcomes of hypothetical interventions rather than just correlating observations, CWMI learns a robust internal representation of physical laws. We demonstrate that our model significantly outperforms state-of-the-art LLMs on zero-shot physical reasoning benchmarks, such as PIQA, and on our newly proposed PhysiCa-Bench dataset. These results validate that inducing an explicit causal world model is a critical and effective step towards building more reliable and generalizable AI systems.
I Introduction
The last few years have witnessed a paradigm shift in artificial intelligence, driven by the advent of Large Language Models (LLMs). Foundational models like BERT [Devlin2019BERT] and the GPT series [Brown2020GPT3, Bubeck2023Sparks] have demonstrated remarkable, often emergent, capabilities across a vast spectrum of natural language tasks. Their proficiency is not limited to text; recent multimodal models such as Gemini [Team2024Gemini] and Flamingo [Alayrac2022Flamingo] have extended this power to encompass vision and other modalities, achieving impressive results in areas like visual dialog and video question answering [Zhang2022Survey]. The capabilities of these models are so extensive that they are often evaluated on large, complex benchmark suites [BIG-Bench-Authors2023] and are increasingly being integrated with external tools to further enhance their functionality [Schick2023Toolformer].
Despite their success in processing and generating human-like language, these models harbor a fundamental limitation: they operate primarily by learning statistical correlations from data, lacking a deep, grounded understanding of the world.This gap becomes particularly apparent when LLMs are confronted with tasks requiring intuitive physical reasoning [Li2023GPT4Physics]. While they can often recite physical laws from their training data, they frequently fail to apply these principles to novel scenarios, a limitation that has been extensively documented [valmeekam2022can]. Advanced prompting strategies such as Chain-of-Thought [Wei2022CoT] and Tree-of-Thoughts [Yao2023Tree] have been proposed to elicit more structured reasoning. However, these methods are essentially techniques for organizing the model’s existing knowledge rather than instilling a genuine, predictive understanding of physical dynamics [Shanahan2022Symbolic]. This deficiency hinders the deployment of LLMs in high-stakes, real-world applications like robotics and autonomous systems, where the ability to predict the consequences of actions is paramount [Ahn2022SayCan, Huang2022InnerMonologue]. The challenge is not merely about recognizing objects, but about understanding their interactions, a problem also explored in audio-visual event analysis .
The core of the issue lies in the distinction between correlation and causation. Current models excel at the former, but true physical intelligence requires the latter. This has led to a growing interest in building models that are more physically grounded [Ding2023LLMReasoning, Li2023Symbolic]. Efforts in this direction include creating embodied multimodal models [driess2023palme], developing new benchmarks for physical and event-based reasoning [Yi2020NSDR, Zellers2019PIQA, Bear2021Physion], and exploring novel architectures for efficient video analysis, such as unified networks for temporal filtering . This is part of a broader trend in creating generative interactive environments from visual data [Team2024Genie]. Concurrently, research in other domains has focused on interpreting the physical world through diverse non-visual signals. WiFi-based sensing, for example, has proven to be a versatile modality. Foundational work established robust human activity recognition by developing methods to mitigate co-channel interference , a crucial step for any real-world interactive system [Park2023Agents]. This reliability is paramount for embodied agents that must ground language in robotic affordances [Ahn2022SayCan]. Building on this, more complex applications like gesture recognition have emerged, where systems must not only be accurate but also handle unknown, open-set gestures to be practical in the real world . The application space for such sensing has expanded dramatically, a trend seen across many forms of human-AI interaction [Wu2023MINT], moving into contactless healthcare with systems that can capture pulmonary function . This extends beyond WiFi, with other RF technologies like RFID-based systems also being explored for identifying fine-grained activities such as writing . Such sensing paradigms enable nuanced tasks that require a form of machine commonsense to interpret [Sap2019Atomic], including unobtrusive emotion recognition . However, the increasing sensitivity of data collected by such pervasive sensing also raises significant security concerns, leading to research into threats like keystroke eavesdropping from commodity devices . These specialized systems, while powerful, often focus on classification rather than building a general, predictive model of the world. The development of such general models is computationally intensive, driving research into distributed learning schemes like federated learning with experience-driven model migration . This is complemented by a focus on algorithmic efficiency through novel architectures [Gu2023Mamba] and advanced techniques like multi-objective model compression .
To bridge this gap, we turn to the concept of ”world models,” which has shown great promise in reinforcement learning for learning predictive models of an environment [Ha2018WorldModels, Hafner2023DreamerV3]. We posit that for an LLM to achieve genuine physical reasoning, it must be endowed with an internal, causal world model. This paper introduces a novel framework, Causal World Model Induction (CWMI), designed to explicitly embed a model of causal physics within an LLM. Our approach moves beyond simple next-token prediction by integrating a specialized Causal Physics Module (CPM) and a new training objective that forces the model to learn cause-and-effect relationships from multimodal data. This is inspired by work on relational inductive biases [Battaglia2018GNNs] and the need for causal perspectives in machine learning [Kiciman2024Causal, Dasgupta2021Grounded]. By training the model to predict the outcomes of hypothetical interventions, we aim to build a more robust and generalizable understanding of physical dynamics, a challenge analogous to suppressing label noise in facial recognition to find the true underlying patterns . Our goal is to create a system that doesn’t just process language, but reasons about the world it describes.
  1. 1.
    We propose a novel framework, Causal World Model Induction (CWMI), that augments a pre-trained LLM with a dedicated Causal Physics Module (CPM) to explicitly learn and represent physical dynamics.
  2. 2.
    We introduce a Causal Intervention Loss ($L_{causal}$), a new training objective that forces the model to learn cause-and-effect relationships by reasoning about counterfactual outcomes, moving beyond standard supervised learning.
  3. 3.
    We develop and will release PhysiCa-Bench, a new multimodal benchmark designed to rigorously evaluate causal physical reasoning in zero-shot settings, with scenarios that probe for genuine understanding over pattern matching.
  4. 4.
    We demonstrate through extensive experiments that our CWMI-enhanced model significantly outperforms state-of-the-art baselines on established benchmarks like PIQA and on our new PhysiCa-Bench, validating that inducing an explicit causal world model is a critical step towards more robust AI.
II Related Work
Our research is situated at the intersection of large language models, world models, and causal reasoning. We review key developments in these areas to contextualize our contributions.
II-A Reasoning in Large Language Models
The capacity for reasoning in LLMs has been a subject of intense investigation. Early work focused on the emergent properties of scaled-up models [Brown2020GPT3]. More recently, research has shifted towards explicit techniques to guide the reasoning process. Chain-of-Thought (CoT) prompting [Wei2022CoT] and its derivatives, such as Tree of Thoughts [Yao2023Tree], have shown that instructing models to ”think step-by-step” can significantly improve performance on arithmetic, commonsense, and symbolic reasoning tasks. Other approaches aim to augment LLMs with external tools or code execution engines, allowing them to offload complex calculations or verify information, as seen in Toolformer [Schick2023Toolformer] and ViperGPT [Li2023ViperGPT]. Frameworks have been proposed to structure the entire comprehension-reasoning-planning process [Zhang2023Igniting]. However, these methods primarily enhance the application of knowledge already stored in the model’s parameters. They do not fundamentally alter the model’s internal representation of the world, which often lacks the physical grounding necessary for robust, real-world interaction [Ding2023LLMReasoning]. This is a critical distinction, as a model can articulate a flawless logical chain yet still base it on a physically implausible premise.
II-B World Models and Physical Prediction
The concept of a ”world model” [Lecun2022Path], an internal, predictive model of an environment, originated largely in the context of reinforcement learning and robotics [Ha2018WorldModels]. Models like DreamerV3 [Hafner2023DreamerV3] learn directly from sensory inputs to create a latent space where planning can occur, enabling agents to master diverse and complex tasks from rich data. This includes understanding and predicting fine-grained micro-actions, a challenge addressed by dedicated benchmarking efforts . In parallel, the computer vision community has developed models for explicit physical prediction from visual data. Physics-informed neural networks (PINNs) embed differential equations directly into the learning process [Raissi2019PINNs]. Other work has focused on creating challenging benchmarks like CLEVRER [Yi2020NSDR] and Physion [Bear2021Physion] to evaluate models on their ability to predict the outcomes of physical events. Recently, generative models like GENIE [Team2024Genie] have shown the ability to create entire interactive environments from a single image, demonstrating a powerful form of implicit world modeling. Our work draws inspiration from these approaches but differs in a key aspect: instead of learning from low-level sensory data or within a constrained simulation, we aim to induce a general, causal world model within a pre-trained LLM using a combination of textual descriptions and video observations. This leverages the vast knowledge already present in the LLM while grounding it in causal physical principles. The development of such complex models relies on advances in underlying neural architectures, such as Mamba [Gu2023Mamba], and efficient learning schemes across heterogeneous networks .
II-C Causality and Multimodal Grounding
A central thesis of our work is that true physical reasoning requires a causal understanding. The field of causal inference provides a formal language for this, but applying it to high-dimensional data like text and video is a significant challenge [Kiciman2024Causal]. Some work has explored using counterfactuals for data augmentation [Goyal2019Counterfactuals] or learning causal relationships in a compositional manner [Dasgupta2021Grounded]. In the multimodal domain, significant effort has been directed at grounding language in perception. This includes pre-training on large-scale video-language datasets [Jin2024VILA, Girdhar2023LAVILA], developing neuro-symbolic architectures that combine perception with explicit reasoning [Mao2019NSCL], and creating embodied agents that learn from interaction [Chen2023VIMA]. The challenge of grounding is also present in other sensing modalities, where researchers aim to prevent attacks on physical-layer authentication . This focus on security and reliability is paramount for embodied agents that must function in noisy, real-world conditions [Ahn2022SayCan]. In a similar vein, others seek to build robust activity recognition systems by developing anti-interference techniques based on phase components . The ability to extract a clean signal from noise is a general problem, with solutions often involving novel neural module communication strategies [Mott2019MIC]. The successful application of these robust techniques enables more advanced, target-oriented sensing in cluttered environments for applications like healthcare . Interpreting these signals to understand high-level human states, such as emotions, requires a form of intuitive physics or theory of mind [Piloto2022Intuitive]. This is exemplified by systems that fuse WiFi and vision for unobtrusive emotion recognition . Our work directly tackles this by using causal interventions as a supervisory signal, forcing the model to learn not just *what* happens, but *why* it happens. This is analogous to efforts in other fields to learn from heterogeneous data sources to a more reliable representation or perform commonsense reasoning over knowledge graphs [Sap2019Atomic].
III Experimental Setup
To validate the effectiveness of our Causal World Model Induction (CWMI) framework, we designed a comprehensive set of experiments. This section details the datasets used for training and evaluation, our implementation and hardware specifications, and the metrics used to measure performance.
III-A Datasets
Our experimental methodology relies on a combination of existing and newly developed datasets to ensure a thorough evaluation of physical reasoning capabilities.
III-A1 Training and Fine-tuning
For the initial pre-training of the Causal Physics Module (CPM), we leveraged a large, curated subset of video-captioning datasets, primarily drawing from Something-Something V2. We filtered this dataset to retain clips that depict clear, unambiguous physical interactions involving a small number of objects (e.g., pushing, dropping, colliding). This resulted in a training set of approximately 200,000 video-text pairs.
For fine-tuning and as a primary evaluation tool, we introduce PhysiCa-Bench, our novel benchmark for Causal Physical Reasoning. PhysiCa-Bench consists of 10,000 high-quality examples. Each example includes: (1) a textual description of a physical scene, (2) a short video clip showing the outcome of the interaction, (3) a multiple-choice question about the direct outcome, and (4) a corresponding counterfactual question that alters one variable of the initial state (e.g., object mass, initial velocity) and asks for the new outcome. This structure is specifically designed to test for genuine causal understanding beyond simple pattern recognition. The data was generated using a physics simulator (PyBullet) and manually validated for quality and realism.
III-A2 Zero-Shot Evaluation
To assess the generalization capability of our model, we perform zero-shot evaluation on the well-established PIQA (Physical Interaction Quality Assessment) benchmark [Zellers2019PIQA]. PIQA is a multiple-choice question-answering dataset that requires commonsense reasoning about how to perform everyday tasks involving physical interaction. It contains 16,113 training examples and 1,838 for the public test set. Crucially, we do not use any of the PIQA training data during the training or fine-tuning of our model to ensure a true zero-shot evaluation setting.
III-B Implementation Details and Hardware
Our framework is built upon a frozen, pre-trained LLM backbone to leverage its extensive world knowledge and language capabilities.
  • Base Model: We use the publicly available Llama 3 8B model as our frozen LLM backbone. Its parameters are not updated during our training process.
  • Causal Physics Module (CPM): The CPM is implemented as a 12-layer Transformer decoder with 8 attention heads and an embedding dimension of 768, resulting in approximately 256 million trainable parameters. It takes the final hidden state of the LLM as input and is trained to predict a structured representation of the final physical state.
  • Training: The model was trained end-to-end (with the LLM frozen) using the AdamW optimizer with a learning rate of $1\times 10^{-5}$ and a cosine decay schedule. We used a global batch size of 128. The hyperparameters for our composite loss function were set to $\alpha=0.5$ and $\beta=1.0$ based on preliminary experiments on a validation set.
  • Hardware: All training and inference were conducted on a high-performance computing cluster equipped with 8 NVIDIA H100 GPUs, each with 80GB of HBM3 memory. Training the final model took approximately 72 hours.
III-C Evaluation Metrics
We employ several metrics to provide a multi-faceted assessment of our model’s performance.
  • Reasoning Accuracy: For the multiple-choice question-answering tasks on PIQA and PhysiCa-Bench, we use standard classification accuracy. This metric measures the percentage of questions for which the model selects the correct answer from the given choices.
  • Future State Prediction Accuracy (FSPA): This metric is specific to our PhysiCa-Bench dataset. The CPM predicts a final state vector containing the positions and velocities of objects. FSPA measures the Mean Squared Error (MSE) between the predicted final state vector and the ground-truth state vector extracted from the simulation, providing a direct measure of the world model’s predictive fidelity.
  • Causal Consistency Score (CCS): Also for PhysiCa-Bench, this metric evaluates the model’s ability to handle counterfactuals. A model is causally consistent on a given example if it correctly answers both the standard question and its corresponding counterfactual question. This is a stricter metric than simple accuracy and is designed to penalize models that guess correctly without consistent reasoning.
IV Methodology
Figure 1: Rich, icon‑enhanced flowchart of the CWMI framework. The diagram illustrates the sequence of operations—from “Input Text” through “LLM Encoding,” “Projection Layer,” “Causal Physics Module,” to final “Output”—with color‑coded modules and visual cues to emphasize each stage’s function.
The central hypothesis of our work is that genuine physical reasoning in artificial intelligence cannot emerge from statistical pattern matching alone. Instead, it requires an internal, predictive, and causal model of the world. Current Large Language Models (LLMs), despite their linguistic prowess [Brown2020GPT3, Team2024Gemini], fundamentally lack this capability, operating as sophisticated stochastic parrots that can describe but not truly understand physical dynamics [valmeekam2022can]. To address this critical gap, we introduce the Causal World Model Induction (CWMI) framework. Our methodology is designed not to simply fine-tune an LLM on physics-related text, but to fundamentally augment it with a new computational module explicitly engineered to simulate cause and effect.
The core philosophy of CWMI is to separate the linguistic, semantic processing capabilities of a pre-trained LLM from the structured, causal simulation of physical dynamics. We leverage the LLM as a powerful ”interface” to the world, capable of parsing complex, unstructured natural language descriptions of physical scenarios into a rich, latent representation. This representation then serves as the initial condition for a dedicated, trainable Causal Physics Module (CPM), which acts as a latent-space physics engine. The entire system is trained end-to-end with a novel composite loss function that combines standard predictive objectives with a crucial Causal Intervention Loss. This loss forces the model to learn not just to predict observed outcomes, but to understand *how* and *why* those outcomes change when specific variables in the initial state are altered. This section provides a comprehensive technical description of the CWMI architecture, the design of the Causal Physics Module, our novel training paradigm, and the inference process for zero-shot physical reasoning.
IV-A System Architecture
The architecture of the CWMI framework is a symbiotic two-component system, as illustrated in Figure 2. It comprises a frozen, pre-trained LLM backbone and our novel, trainable Causal Physics Module (CPM). This decoupled design is intentional: it allows us to harness the vast semantic knowledge and linguistic fluency of a state-of-the-art LLM without the catastrophic forgetting or computational expense associated with full fine-tuning, while concentrating the learning of physical dynamics within a specialized, efficient module.
Figure 2: The overall architecture of the Causal World Model Induction (CWMI) framework. An input text describing a physical scene is processed by the frozen LLM backbone. The LLM’s final hidden state, which encodes a rich semantic representation of the scene, is used to initialize the state of the trainable Causal Physics Module (CPM). The CPM, acting as a latent physics simulator, predicts the final state of the system. The entire model is trained with a composite loss function, including a key Causal Intervention Loss derived from counterfactual pairs.
The information flow through the system at training time proceeds as follows:
  1. 1.
    Input Processing: The system receives a natural language text prompt $\mathcal{T}$ describing an initial physical scene. For example, $\mathcal{T}$ could be ”A small red rubber ball is on a steep metal ramp above a wooden floor.”
  2. 2.
    LLM Encoding: The prompt $\mathcal{T}$ is tokenized and fed into the frozen LLM backbone (e.g., Llama 3 8B). The LLM processes the text through its multiple transformer layers. We extract the final hidden state vector, $h_{\mathcal{T}}\in\mathbb{R}^{d_{llm}}$, corresponding to the last token of the input sequence. This vector serves as a compressed, semantically rich representation of the entire scene, capturing the objects, their properties (e.g., ”small,” ”red,” ”rubber”), and their spatial relationships (”on,” ”above”).
  3. 3.
    State Initialization: The LLM’s hidden state $h_{\mathcal{T}}$ is projected into the latent space of the Causal Physics Module via a linear transformation, forming the initial state of the physical simulation, $S_{0}\in\mathbb{R}^{d_{cpm}}$. This projection acts as a bridge, translating from the LLM’s semantic space to the CPM’s physics-oriented state space.
    $S_{0}=W_{proj}h_{\mathcal{T}}+b_{proj}$ (1)
    where $W_{proj}\in\mathbb{R}^{d_{cpm}\times d_{llm}}$ and $b_{proj}\in\mathbb{R}^{d_{cpm}}$ are trainable parameters.
  4. 4.
    Causal Simulation: The CPM takes the initial state $S_{0}$ and iteratively, or in a single step, computes a predicted final state, $\hat{S}_{final}\in\mathbb{R}^{d_{cpm}}$. This process is designed to model the temporal evolution of the physical system, governed by the causal laws learned during training.
  5. 5.
    Loss Computation: The predicted final state $\hat{S}_{final}$ is compared against a ground-truth final state $S_{final}$, which is derived from the ground-truth outcome (e.g., from a video in our PhysiCa-Bench dataset). This comparison forms the basis of our predictive and causal loss functions, which will be detailed in Section 3.3. The gradients are computed and used to update only the parameters of the CPM and the projection layer.
This architectural choice ensures a clear division of labor. The LLM is the expert linguist and knowledge extractor, while the CPM is the apprentice physicist, learning the laws of motion and interaction from data.
IV-B The Causal Physics Module (CPM)
The Causal Physics Module is the heart of our framework. It is not merely a classifier or a regressor; it is designed to function as a differentiable physics engine operating in a learned latent space. Its purpose is to take an initial state representation and predict a future state by simulating the underlying physical processes.
IV-B1 State Representation
A critical design choice is how to represent a physical scene within the CPM’s latent space. Inspired by relational inductive biases in graph networks [Battaglia2018GNNs] and object-centric learning, our state vector $S\in\mathbb{R}^{d_{cpm}}$ is structured to implicitly represent the key entities and their properties. While not explicitly factored into distinct object slots in the current implementation, the dimensionality and training objective encourage the model to learn a disentangled representation where different dimensions correspond to properties like position, velocity, mass, and material of the primary objects in the scene. For a scene with up to $N$ objects, each with $P$ properties, the state dimension $d_{cpm}$ would be conceptually related to $N\times P$. The LLM-to-CPM projection layer is tasked with the crucial role of extracting this structured information from the LLM’s dense, holistic embedding.
IV-B2 CPM Architecture
The CPM itself is implemented as a Transformer-based architecture, chosen for its proven ability to model complex relationships and transformations within sequential data. Specifically, we use a Transformer decoder stack with $L=12$ layers. The initial state vector $S_{0}$ is treated as the first element in a sequence. The module uses self-attention mechanisms to model the interactions between the different components of the state vector.
The operation within a single CPM layer can be described as:
$S^{\prime}$ $=\text{LayerNorm}(S+\text{MultiHeadSelfAttention}(S))$ (2)
$S_{new}$ $=\text{LayerNorm}(S^{\prime}+\text{FeedForwardNetwork}(S^{\prime}))$ (3)
This process is repeated for $L$ layers. The multi-head self-attention mechanism allows the module to learn a relational structure, where, for instance, the future state of one object (e.g., a ball) is computed by attending to the state of other objects it might interact with (e.g., a ramp, the floor). The feed-forward network then applies a non-linear transformation, modeling the complex physics of the interaction. The output of the final layer is the predicted final state, $\hat{S}_{final}$. This single-pass, deep transformation is designed to approximate the integrated result of a continuous physical evolution over a short time horizon.
IV-C Causal Induction Training Objective
The most significant methodological innovation of our work lies in the training objective. A model trained solely to predict observed outcomes can achieve low error by becoming a master of correlation, but it may not learn the underlying causal mechanisms. For example, in a dataset where heavy objects are always large, a model might incorrectly learn that ”size” causes faster falling, rather than ”mass.” To overcome this, we introduce a composite loss function that explicitly forces the model to reason about causality through intervention.
The total loss, $L_{total}$, is a weighted sum of a standard state prediction loss and our novel Causal Intervention Loss:
$L_{total}=\alpha L_{pred}+\beta L_{causal}$ (4)
We optionally include a standard language modeling loss, $L_{LM}$, if we were to fine-tune the LLM, but in our primary framework with a frozen LLM, this term is omitted. The hyperparameters $\alpha$ and $\beta$ control the relative importance of each objective.
IV-C1 State Prediction Loss ($L_{pred}$)
The State Prediction Loss is a supervised objective that anchors the model’s predictions to reality. It measures the discrepancy between the CPM’s predicted final state $\hat{S}_{final}$ and the ground-truth final state $S_{final}$. The ground-truth state $S_{final}$ is obtained by encoding the ground-truth outcome (e.g., the final frame of a video) using a pre-trained, non-trainable visual encoder (e.g., a frozen ViT) whose output is projected into the CPM’s state space. This loss ensures the CPM learns to produce physically plausible outcomes that match the observed data. We use the Mean Squared Error (MSE) for its simplicity and effectiveness:
$L_{pred}=\frac{1}{d_{cpm}}\sum_{i=1}^{d_{cpm}}(\hat{S}_{final,i}-S_{final,i})^{2}$ (5)
This loss component teaches the model ”what” happens in a typical scenario.
IV-C2 Causal Intervention Loss ($L_{causal}$)
The Causal Intervention Loss is the key to inducing a causal world model. It moves beyond passive observation and forces the model to reason about the consequences of actions or changes—the essence of causal understanding. This loss is computed using pairs of factual and counterfactual scenarios from our PhysiCa-Bench dataset.
A factual scenario consists of an initial description $\mathcal{T}$ and its observed outcome, leading to a final state $S_{final}$. A counterfactual scenario consists of a slightly modified description $\mathcal{T}^{\prime}$ (e.g., ”the rubber ball is now a heavy steel ball”) and its corresponding outcome, leading to a different final state $S^{\prime}_{final}$.
The Causal Intervention Loss is designed to penalize the model if its prediction for the counterfactual scenario is inconsistent with the actual effect of the intervention. Specifically, we want the change in the model’s predicted state to match the change in the ground-truth state.
Let $\hat{S}_{final}$ be the model’s prediction for the factual input $\mathcal{T}$, and $\hat{S}^{\prime}_{final}$ be the prediction for the counterfactual input $\mathcal{T}^{\prime}$. The predicted change due to the intervention is the vector difference $\Delta\hat{S}=\hat{S}^{\prime}_{final}-\hat{S}_{final}$. Similarly, the true change is $\Delta S=S^{\prime}_{final}-S_{final}$. The Causal Intervention Loss measures the difference between these two ”change vectors”:
$L_{causal}=\frac{1}{d_{cpm}}\sum_{i=1}^{d_{cpm}}((\hat{S}^{\prime}_{final,i}-\hat{S}_{final,i})-(S^{\prime}_{final,i}-S_{final,i}))^{2}$ (6)
Physical Meaning: This loss function has a profound physical and causal interpretation. It does not reward the model for correctly predicting the absolute final states $\hat{S}_{final}$ or $\hat{S}^{\prime}_{final}$—that is the job of $L_{pred}$. Instead, $L_{causal}$ focuses exclusively on whether the model correctly predicts the *effect* of the intervention. For example, if changing a ball from rubber to steel (the intervention) causes its final velocity to increase by 5 m/s (the effect), this loss function pushes the model to learn precisely that relationship. It forces the model to isolate the causal link between a specific change in an initial condition and the resulting change in the outcome. By training on a diverse set of such interventions (changing mass, friction, initial velocity, etc.), the CPM learns a robust model of physics that is less susceptible to spurious correlations. This approach is a practical implementation of the ideas from structural causal models [Kiciman2024Causal] and counterfactual reasoning [Goyal2019Counterfactuals] tailored for deep learning models.
IV-D Zero-Shot Inference for Physical QA
Once the CWMI framework is trained, it can be used to answer new, unseen physical reasoning questions in a zero-shot manner. The inference process leverages both the CPM’s simulation capabilities and the LLM’s language generation fluency.
Consider a multiple-choice question from a benchmark like PIQA [Zellers2019PIQA]. The input consists of a premise (e.g., ”To dry a shirt, you hang it on a clothesline.”) and two possible solutions (e.g., Sol 1: ”The sun and wind evaporate the water.”; Sol 2: ”The shirt absorbs sunlight, making it heavier.”).
The inference procedure is as follows:
  1. 1.
    Scenario Formulation: For each possible solution, we create a complete physical scenario by concatenating the premise and the solution.
    • Scenario 1: ”To dry a shirt, you hang it on a clothesline. The sun and wind evaporate the water.”
    • Scenario 2: ”To dry a shirt, you hang it on a clothesline. The shirt absorbs sunlight, making it heavier.”
  2. 2.
    Plausibility Scoring: Each scenario string is passed through the CWMI model to obtain a ”plausibility score.” This score is derived from the internal consistency between the LLM’s understanding and the CPM’s simulation. We compute the predicted final state $\hat{S}_{final}$ using the CPM. The plausibility score is defined as the negative prediction error (e.g., negative MSE) between the LLM’s initial representation of the full scenario and the outcome simulated by its own physics module. A more physically plausible scenario will result in a smaller internal prediction error, and thus a higher score.
    A simpler and more direct method, which we use in our experiments, is to compute a pseudo-log-likelihood score. We take the final state predicted by the CPM, $\hat{S}_{final}$, project it back into the LLM’s embedding space, and use it to condition the LLM backbone to generate a textual description of the outcome. The score for a given choice is the conditional probability of generating that choice’s text given the premise.
    $\text{Score}(\text{Sol}_{j}|\text{Premise})=P_{LLM}(\text{Sol}_{j}|\text{Premise},\text{CPM}(\text{Premise}))$ (7)
  3. 3.
    Answer Selection: The model selects the solution corresponding to the scenario that received the highest plausibility score.
This inference mechanism allows the model to ”test” different hypotheses. It uses its internal, learned physics engine (the CPM) to simulate the consequences of each potential answer and selects the one that is most consistent with its learned model of the world. This is a significant departure from standard LLMs, which would select an answer based purely on statistical correlations in their training text, and is the key to our model’s enhanced zero-shot reasoning capabilities.
V Results and Discussion
The primary objective of our empirical evaluation is to rigorously validate the central hypothesis of this paper: that explicitly inducing a causal world model within a Large Language Model is a necessary and effective strategy for achieving robust physical reasoning. To this end, we conducted a series of experiments designed to assess the performance of our Causal World Model Induction (CWMI) framework against state-of-the-art baselines, dissect the contributions of its core components, and gain qualitative insights into its operational behavior. This section presents a comprehensive analysis of our findings. We begin by reporting quantitative results on two distinct benchmarks—the established PIQA dataset and our novel, causality-focused PhysiCa-Bench. We then delve into a series of ablation studies to isolate the impact of our key methodological innovations. Finally, we discuss insights gleaned from model scaling and qualitative error analysis, contextualizing our achievements within the broader landscape of artificial intelligence research.
V-A Quantitative Analysis of Zero-Shot Reasoning
A key desideratum for any advanced reasoning system is the ability to generalize to new, unseen problems without task-specific fine-tuning. We therefore evaluated our CWMI model in a challenging zero-shot setting on both a commonsense physical reasoning task and our targeted causal reasoning benchmark.
V-A1 Performance on the PIQA Benchmark
We first assessed our model on the Physical Interaction Quality Assessment (PIQA) benchmark [Zellers2019PIQA], which requires commonsense understanding of everyday physical situations. The results, summarized in Table I, demonstrate a clear and significant advantage for our approach. Our CWMI model achieves a zero-shot accuracy of 89.4%, outperforming the base Llama 3 8B model by a substantial margin of 15.2 percentage points. More notably, it also surpasses the reported performance of much larger, proprietary models like GPT-4, which, despite their vast parameter counts and training data, still struggle with the nuances of physical common sense [Li2023GPT4Physics, Bubeck2023Sparks].
Figure 3: Zero-Shot Reasoning Accuracy on PIQA.
TABLE I: Zero-Shot Reasoning Accuracy on the PIQA Test Set. Our CWMI framework significantly outperforms both its base model and larger, general-purpose LLMs, highlighting the benefit of an explicit world model for physical commonsense.
Model Accuracy (%)
Random Chance 50.0
Llama 3 8B (Base Model) 74.2
GPT-4 (Reported in [Li2023GPT4Physics]) 85.5
CWMI (Ours) 89.4
The performance gap, visualized in the bar chart , is telling. Standard LLMs approach physical reasoning as a language modeling problem; they select the answer that is most statistically probable given their training corpus [Shanahan2022Symbolic]. This makes them vulnerable to ”fluent fallacies”—answers that are linguistically plausible but physically nonsensical. Our CWMI framework, by contrast, approaches the problem as a simulation task. As described in our methodology (Section IV), it uses its Causal Physics Module (CPM) to internally simulate the outcome of each potential solution and selects the one that is most physically coherent. This ability to perform latent-space simulation provides a powerful inductive bias that is absent in standard architectures. While advanced prompting techniques like Chain-of-Thought [Wei2022CoT] can elicit better reasoning from base models, they are fundamentally post-hoc strategies that attempt to structure the model’s existing, flawed knowledge. Our work, in contrast, directly addresses the flaw in the underlying knowledge itself by embedding a more robust, causal model of the world. This result validates that for tasks grounded in the physical world, a specialized, model-based reasoning component is superior to relying on the emergent, and often brittle, capabilities of a pure language model.
V-A2 Performance on the PhysiCa-Bench Benchmark
While PIQA is an excellent measure of commonsense reasoning, we designed PhysiCa-Bench to more directly and rigorously probe the causal understanding of a model. This benchmark evaluates models not only on their ability to predict outcomes but also on their ability to correctly predict the effect of specific interventions. The results, presented in Table II, reveal the profound limitations of existing models and the unique strengths of our framework.
TABLE II: Performance on the PhysiCa-Bench Test Set. Our model excels across all metrics, particularly the strict Causal Consistency Score (CCS), demonstrating a true understanding of cause and effect that baseline models lack.
Model Accuracy (%) FSPA (MSE $\downarrow$) CCS (%)
Llama 3 8B 61.7 N/A 12.3
GPT-4 78.5 N/A 21.9
CWMI (Ours) 94.1 0.08 87.6
On standard reasoning accuracy, CWMI achieves an impressive 94.1%, far surpassing the baselines. This is expected, as the training objective of our model is closely aligned with the nature of the task. However, the more insightful metrics are the Future State Prediction Accuracy (FSPA) and the Causal Consistency Score (CCS).
The FSPA directly measures the fidelity of the CPM as a physics simulator by calculating the Mean Squared Error between its predicted final state and the ground-truth state from the simulation. Our model achieves a low FSPA of 0.08, indicating that its internal world model generates accurate, quantitative predictions. Baseline LLMs cannot be evaluated on this metric as they lack the architectural components to produce a structured state prediction, a key differentiator of our approach. This capacity for state prediction is a foundational element of classic world models in robotics and reinforcement learning [Ha2018WorldModels, Hafner2023DreamerV3], and our work successfully integrates this principle into the LLM paradigm.
The most compelling evidence for our framework’s success comes from the Causal Consistency Score (CCS). This strict metric requires the model to correctly answer both a factual question and its corresponding counterfactual counterpart. A high CCS is difficult to achieve through superficial pattern matching; it requires the model to understand the specific causal link between the variable that was changed (the intervention) and the change in the outcome. Our CWMI model achieves a remarkable CCS of 87.6%. In stark contrast, Llama 3 and GPT-4 score only 12.3% and 21.9%, respectively. Their scores, while better than random chance, suggest they are answering the questions independently, with no consistent underlying causal model to link the factual and counterfactual pairs. They may correctly guess that a steel ball falls faster than a rubber one, but they fail to do so with the consistency that would imply genuine understanding. This finding strongly supports our core thesis: the interventional training paradigm described in Section IV (Eq. 6) is exceptionally effective at forcing a model to move beyond correlation and learn the causal structure of the physical world. This aligns with broader calls in the AI community for a more central role for causality in building robust and generalizable intelligence [Kiciman2024Causal, Dasgupta2021Grounded].
V-B Ablation Studies: Deconstructing the CWMI Framework
To understand the source of CWMI’s performance and validate our specific design choices, we conducted a series of ablation studies on PhysiCa-Bench. We systematically removed or altered key components of our framework and measured the impact on performance.
TABLE III: Ablation Study of the CWMI Framework on PhysiCa-Bench…
Model Variant Accuracy (%) FSPA (MSE $\downarrow$) CCS (%)
Full CWMI Model 94.1 0.08 87.6
CWMI w/o $L_{causal}$ 85.3 0.11 43.2
CWMI w/o $L_{pred}$ 79.8 0.45 51.5
LLM Fine-tuning (No CPM) 68.1 N/A 19.8
V-B1 The Critical Role of the Causal Intervention Loss
The most crucial component of our framework is the Causal Intervention Loss ($L_{causal}$). To quantify its impact, we trained a variant of our model using only the State Prediction Loss ($L_{pred}$). This model, labeled ”CWMI w/o $L_{causal}$”, was trained to predict the final state from a given input but never explicitly trained on counterfactual pairs.
V-B2 The Grounding Function of the State Prediction Loss
We also trained a variant without the State Prediction Loss ($L_{pred}$), relying only on the Causal Intervention Loss. The goal was to see if a model could learn physics purely by observing the *effects* of changes. The results for ”CWMI w/o $L_{pred}$” show that this is not an effective strategy. While the CCS remains higher than the baseline LLMs, it is significantly lower than the full model. More importantly, the FSPA balloons to 0.45, indicating that the model’s absolute predictions are no longer well-grounded in physical reality.
This outcome is intuitive. The $L_{causal}$ term (Eq. 6) is a differential loss; it cares only about the *difference* between the factual and counterfactual outcomes. Without the $L_{pred}$ term to anchor the predictions to the ground truth, the model is free to learn the correct ”change dynamics” within a distorted, physically implausible latent space. For example, it might correctly learn that gravity causes downward acceleration, but its simulation could take place in a world where the ground plane is meters away from its actual position. The $L_{pred}$ term acts as a crucial grounding mechanism, ensuring that the world model is not only causally consistent but also observationally accurate. The two loss terms work in synergy: $L_{pred}$ teaches the model ”what the world looks like,” while $L_{causal}$ teaches it ”how the world works.”
V-B3 Validating the Architectural Separation
Finally, to justify our two-component architecture, we conducted an experiment where we discarded the CPM entirely and instead fine-tuned the end layers of the Llama 3 LLM directly on the physical reasoning tasks. This ”LLM Fine-tuning (No CPM)” approach is representative of standard domain adaptation techniques. Its performance is poor across the board, barely improving upon the zero-shot GPT-4 baseline and showing almost no causal consistency. This result strongly validates our architectural decision to separate linguistic parsing from physical simulation. A standard Transformer architecture, while a universal approximator in theory, lacks the specific inductive biases for modeling physical dynamics efficiently. Our CPM, while also Transformer-based, is specialized for this role. It operates on a structured state representation and is trained with a physics-oriented objective, making it far more effective and parameter-efficient for this class of problems. This finding echoes the principles of modularity and relational inductive biases [Battaglia2018GNNs, Mott2019MIC], suggesting that building complex AI systems may require composing specialized modules rather than relying on a single, monolithic model for all tasks.
V-C Model Scaling and Qualitative Insights
Beyond aggregate metrics, a deeper understanding of our model’s behavior can be gained by analyzing how its performance scales and by examining its successes and failures on specific examples.
V-C1 The Effect of Causal Physics Module Capacity
Figure 4: CWMI Performance vs CPM Capacity: Causal Consistency Score
This trend is highly informative. The initial sharp increase in performance suggests that a certain model capacity is required to capture the fundamental physical laws present in our dataset. As the model grows, it becomes more adept at modeling complex interactions and second-order effects. The eventual plateau indicates that, for the complexity of physics represented in PhysiCa-Bench (primarily rigid-body dynamics), our 256M-parameter CPM is near-optimal. This scaling analysis provides confidence in our model design and suggests a clear path for future work: to tackle more complex physical domains like fluid or non-rigid dynamics, a correspondingly larger and perhaps architecturally different CPM will be required. This mirrors the scaling laws that have been so fundamental to the progress of LLMs themselves [Brown2020GPT3].
V-C2 Qualitative Analysis: When It Works and When It Fails
A qualitative examination of the model’s predictions provides the most intuitive feel for its capabilities. On a typical PhysiCa-Bench example—e.g., ”A hockey puck slides across a wooden surface and stops. What happens if the surface is changed to ice?”—baseline LLMs often provide generic or incorrect answers. GPT-4 might correctly state the puck ”slides farther,” but it cannot provide a consistent quantitative prediction. Our CWMI model, however, not only selects the correct qualitative answer but its internal CPM predicts a final state vector with a significantly lower friction coefficient and a correspondingly higher final velocity.
The model’s successes are rooted in its ability to ground abstract concepts in its simulated physics. When it processes ”ice,” the LLM’s initial embedding activates dimensions in the CPM’s state space that correspond to a low coefficient of friction, which the CPM has learned will lead to less deceleration. This seamless link between semantic understanding and physical simulation is the hallmark of our framework.
However, the model is not infallible. Our error analysis revealed several recurring failure modes. The current model struggles with:
  1. 1.
    Complex Multi-Object Interactions: While it can handle two- or three-body collisions, scenarios involving many simultaneous interactions (e.g., a pile of falling blocks) lead to chaotic and inaccurate predictions. Its state representation may not have sufficient capacity to track many objects distinctly.
  2. 2.
    Non-Rigid and Fluid Dynamics: The model has been trained on rigid-body physics. When presented with questions involving cloth, liquids, or deformable objects, its predictions are no better than a standard LLM’s guess.
  3. 3.
    Subtle or Implicit Physical Properties: The model fails if a key physical property is not explicitly stated or strongly implied in the text. For example, it does not implicitly understand that a ”feather” has high air resistance unless prompted.
These limitations are not surprising; they define the boundaries of our current contribution and illuminate the path forward. Addressing them will likely require architectural enhancements to the CPM, such as adopting graph neural networks for multi-object tracking [Battaglia2018GNNs], incorporating physics-informed neural network techniques for continuous fields [Raissi2019PINNs], or integrating more sophisticated 3D visual representations [Lerer2019Pytorch3D]. Furthermore, creating more robust systems that can handle noisy or incomplete descriptions is a shared challenge across many AI domains, from facial expression recognition from heterogeneous sources to building reliable WiFi-based sensing systems in the face of interference [9613773]. Our work represents a foundational step in building a robust reasoning engine, but true physical intelligence will require continued progress on all these fronts.